# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 6 Case study: the EXTRA algorithm
  Choose step-size $\alpha>0$, matrices $W,\leavevmode\nobreak\ \tilde{W}\in\mathbb{R}^{n\times n}$ and pick any $x_{i}^{0}\in\mathbb{R}^{d}$ ($\forall i$);
@@
[…]
@@
[…]
@@
Bounds (120) and (121) directly follows from [14, Corollary 2]. ∎
@@ 6.1 Performance of the worst agent
Figure 1: Comparison of the average functional error $E_{f}$ (122) and the worst functional error $E_{f,\mathrm{worst}}$ (123) for $t=15$ iteration of the EXTRA algorithm and their evolution with the number of agents $n$ in the system. The plot shows (i) the constant theoretical bound on $E_{f}$ (120), in black, (ii) the PEP bounds on the function error of the worst agent $E_{f,\mathrm{worst}}$, scaling sublinearly with $n$, in red and (iii) the PEP bounds for the function error of the average iterate $E_{f}$, in blue. The constant step-size $\alpha^{*}$ has been optimized for this performance criterion and leads to a bound constant with $n$ (plain blue line). Step-sizes can also be optimized with respect to $E_{f,\mathrm{worst}}(15)$, leading to smaller step-sizes $\alpha^{*}_{\mathrm{w}}(n)$ decreasing as $\frac{1}{\sqrt{n}}$. They improve the guarantee for $E_{f,\mathrm{worst}}$ and its scaling with $n$ but they deteriorate the corresponding guarantee for $E_{f}$ (see dashed lines). In this plot, the range of eigenvalues for the averaging matrix is [0.5,0.5]delimited-[]05.05[-0.5,0.5][ - 0.5,0.5 ] and the local functions are 1-smooth and 0.1-strongly convex.
@@
The maximization of such objectives ensures that agent 1 will be the one with the largest error (on $f$ or $x$). Indeed, if another agent $j$ has a larger error, then the permutation of variables of agents $1$ and $j$ in the PEP solution would lead to another feasible PEP solution with a larger objective. Agent 1 will thus receive a specific role in the PEP solution and is not equivalent to the $n-1$ others. We can formulate a compact PEP using two equivalence classes of agents
@@
Firstly, Figure 1 shows that the PEP bound for the average performance $E_{f}$ is well constant over $n$, as predicted by Theorem 7, and outperforms the theoretical bound from [14, Corollary 2], which requires 2750 iterations to guarantee the same error as the PEP bounds after 15 iterations. The theoretical bound is valid for a step-size $\alpha=\frac{1}{4L}$, while the PEP bound is shown for an optimized step-size $\alpha^{*}=\frac{0.78}{L}$, which minimizes the bound. Since the bound is constant over $n$, the value of the optimized step-size can be used for any system size, however, it may depend on other settings, such as the range of eigenvalues for $W$: $[\lambda^{-},\lambda^{+}]$.
There is currently no theoretical guarantee for the worst agent functions error $E_{f,\mathrm{worst}}$ of EXTRA in the literature, but we can analyze it with our compact PEP formulation. Figure 1 shows that the PEP bound for the worst agent $E_{f,\mathrm{worst}}$ (in red) is scaling sublinearly with the number of agents $n$. Using the constant step-size $\alpha^{*}=\frac{0.78}{L}$ (plain red line), we obtain a logarithmic slope of 0.920920.920.92. The step-size can also be tuned with respect to $E_{f,\mathrm{worst}}$, and would then have a different value for each value of $n$, denoted $\alpha^{*}_{\mathrm{w}}(n)$. Using $\alpha^{*}_{\mathrm{w}}(n)$ (dashed red line), we obtain a logarithmic slope of 0.670670.670.67, which means the scaling with $n$ is less important when using suitable step-sizes. In our case, the step-sizes $\alpha^{*}_{\mathrm{w}}(n)$ are diminishing in $\frac{1}{\sqrt{n}}$. However, these smaller step-sizes deteriorate the average performance $E_{f}$ (dashed blue line).
@@
There is currently no theoretical guarantee for the worst agent iterates error $E_{x,\mathrm{worst}}$ of EXTRA in the literature, but one could derive222using the fact that $\max_{i\in\mathcal{V}}\|x_{i}^{t}-x^{*}\|^{2}\leq\sum_{i=1}^{n}\|x_{i}^{t}-x^{*}\|^{2}$ a conservative bound on $E_{x,\mathrm{worst}}$ from (121), scaling linearly with the number of agents $n$. Using our compact PEP formulation, we can analyze such performance metric accurately and show that the performance of the worst agent $E_{x,\mathrm{worst}}$ for EXTRA actually scales sublinearly with $n$, as shown in Figure 2. Figure 2 shows, in particular, PEP-based bounds for $E_{x}$ and $E_{x,\mathrm{worst}}$. The observations are similar to those of Figure 1. The PEP bound for the average performance $E_{x}$ is constant over $n$, as predicted by Theorem 7 and the corresponding optimized step-size is $\alpha^{*}=\frac{0.78}{L}$. Moreover, the PEP bound for the worst agent $E_{x,\mathrm{worst}}$ is scaling sublinearly with the number of agents $n$. Using the constant step-size $\alpha^{*}=\frac{0.78}{L}$ (plain red line), we obtain a logarithmic slope of 0.820820.820.82. As previously, the step-size can also be tuned with respect to $E_{x,\mathrm{worst}}$, and would then have a different value for each value of $n$, denoted $\alpha^{*}_{\mathrm{w}}(n)$. Using the optimal diminishing step-sizes $\alpha^{*}_{\mathrm{w}}(n)$ (dashed red line), we obtain a logarithmic slope of 0.600600.600.60, which means the scaling with $n$ is less important when using suitable step-sizes. These smaller step-sizes only slightly deteriorate the average performance $E_{x}$ (dashed blue line).
@@ 6.2 The 80-th percentile of the agent performance
Figure 2: Comparison of the average iterates error $E_{x}$ (122), the worst iterates error $E_{x,\mathrm{worst}}$ (123) and the 80-th percentile iterates error $E_{x,80}$ (124) for $t=15$ iterations of the EXTRA algorithm and their evolution with the number of agents $n$ in the system. The plot shows (i) the PEP bounds on the iterates error of the worst agent $E_{x,\mathrm{worst}}$, scaling sublinearly with $n$, in red, (ii) the PEP bound on the 80-th percentile iterates error $E_{x,80}$, in green, and (iii) the PEP bounds for the average iterates error $E_{x}$, in blue. The constant step-size $\alpha^{*}$ has been hand-tuned for this performance criterion and leads to a bound constant with $n$ (plain blue line). Step-sizes can also be hand-tuned with respect to $E_{f\mathrm{worst}}(15)$, leading to smaller step-sizes $\alpha^{*}_{\mathrm{w}}(n)$. They improve the guarantee for $E_{f,\mathrm{worst}}$ and its scaling with $n$ without deteriorating too much the corresponding guarantee for $E_{f}$ (see dashed lines). In this plot, the range of eigenvalues for the averaging matrix is [0.5,0.5]delimited-[]05.05[-0.5,0.5][ - 0.5,0.5 ] and the local functions are 1-smooth and 0.1-strongly convex.
In some cases, the performance of the worst agent may not be the most appropriate performance measure. When the network is very large, the performance of the worst agent can be terribly bad, while the average performance is good. Using our PEP approach, we can consider other performance measures that were so far completely out of reach of theoretical analysis. Inspired by statistical approaches, we propose here to analyze the k-th percentile in the distribution of the individual performance of each agent. In this experiment we have chosen the 80-th percentile, that is the performance of the worst agent, after excluding the 20% worst agents. In other words, it measures the error at or below which 80% of the agents fall in the worst-case scenario. The performance of the worst agent, analyzed in the previous subsection, can be seen as the 100-th percentile of the agent performance and can be analyzed via PEP using two equivalence classes of agents. When considering smaller percentiles, such as the 80-th, we can formulate a compact PEP using three equivalence classes of agents:
@@
where we assume $n$ can be divided by 5. The class $\mathcal{V}_{1}$ contains the 20% worst agents to exclude, the class $\mathcal{V}_{2}$ contains the agent for which the individual performance will provide the 80-th percentile, and $\mathcal{V}_{3}$ contains all the other agents. The PEP will therefore maximize the performance of agent $i\in\mathcal{V}_{2}$ and imposes that the performance of each agent $j\in\mathcal{V}_{j}$ is larger:
$E_{x,80}(t)=\underset{\{f_{A}^{u},G_{A}^{u},G_{B}^{u}G_{E}^{uv}\}_{u,v=1,2,3}}{\mathrm{maximum}}$ $\leavevmode\nobreak\ \|x_{i}^{t}-x^{*}\|^{2}\quad\text{ with $i\in\mathcal{V}_{2}$}$ (124)
[…]
In problem (124), the squared norms $\leavevmode\nobreak\ \|x_{i}^{t}-x^{*}\|^{2}$ can be written as $(e_{x^{t}}-e_{x^{*}})^{T}G_{A}^{u}(e_{x^{t}}-e_{x^{*}})$ with $u$ chosen according to the class $\mathcal{V}_{u}$ containing agent $i$. Details about the explicit construction of compact PEP formulations are given in Appendix A.
@@
[…]
@@
[…]
Figure 3: Evolution with k of the k-th percentile of the agent performance for $t$ iterations of EXTRA, when the number of agents tends to infinity ($n\to\infty$). Different total numbers of iterations $t$ are compared. In this plot, the step-size is $\alpha=0.78$, the range of eigenvalues for the averaging matrix is [0.5,0.5]delimited-[]05.05[-0.5,0.5][ - 0.5,0.5 ] and the local functions are 1-smooth and 0.1-strongly convex.
@@ 6.3 Performance under local functions heterogeneity
In many applications, the agents hold local functions sharing general properties, e.g. convexity or smoothness, but it is unlikely that all the local functions have exactly the characterization for these properties. For example, the local functions can be $\mu_{i}$-strongly convex and $L_{i}$-smooth, but with different parameters $\mu_{i}$ and $L_{i}$. In general, the performance is then computed by considering that all the local functions have the worst parameters:
@@
In this experiment, we analyze the gain in the performance guarantee when instead of considering $f_{i}\in\mathcal{F}_{\mu,L}$ for all $i$, we consider two equivalence classes of agents $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ such that
@@
Figure 4: Evolution with the number of agents $n$ of the worst-case average distance to optimum $E_{x}$ for 15 iterations of EXTRA with two equivalence classes of agents $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ using each a different condition number for their function class. The agents in $\mathcal{V}_{1}$ hold 1-smooth and 0.01-strongly convex local functions ($\kappa_{1}=100$) and the agents in $\mathcal{V}_{2}$ hold 1-smooth and 0.1-strongly convex local functions ($\kappa_{2}=10$). The guarantees are independent of the total number of agents $n$ and only depend on the proportion of agents in each class. In this experiment, the range of eigenvalues for the averaging matrix is [0.5,0.5]delimited-[]05.05[-0.5,0.5][ - 0.5,0.5 ].
Figure 4 shows the situation where $L=1$, $\mu_{1}=0.01$ and $\mu_{2}=0.1$, for different sizes of classes $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. We observe that the guarantees for EXTRA only depend on the relative composition of each class $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ but not on the total number of agents $n$. This observation has been confirmed by solving the compact PEPs for $n$ tending to $\infty$, as described at the end of Section 5. In Figure 4, when $20\%$ of the agents pass from the class $\mathcal{V}_{1}$ to the class $\mathcal{V}_{2}$, the worst-case guarantee is improved by a constant factor 0.9090.90.9. This improving factor can be linked to the worst-case guarantees obtained with uniform value for $\kappa$ in the local functions ($\kappa=100$ or $\kappa=10$ for all the agents) and motivates the following conjecture:
@@
[…]
@@
[…]

### v2

@@ 6 Case study : the EXTRA algorithm
 Choose step-size $\alpha>0$, matrices $W,\penalty 10000\ \tilde{W}\in\mathbb{R}^{n\times n}$ and pick any $x_{i}^{0}\in\mathbb{R}^{d}$ ($\forall i$) ;
@@
    The averaging matrix $W\in\mathcal{W}_{\text{\footnotesize$(-\lambda,\lambda)$}}$, with $\lambda\in[0,1)$, and $\tilde{W}=\frac{W+I}{2}$ ;
@@
    $\|x_{i}^{0}-x^{*}\|^{2}\leq R_{1}\qquad\text{and}\qquad\|\nabla f_{i}(x^{*})\|^{2}\leq R_{2}\qquad\text{for all $i=1,\dots,n$\penalty 10000\thinspace;}$
@@
Bounds (117) and (118) directly follows from [15, Corollary 2]. ∎
@@ 6.1 Performance of the worst agent
Figure 1: Comparison of the average functional error $E_{f}$ (119) and the worst functional error $E_{f,\mathrm{worst}}$ (120) for $t=15$ iterations of the EXTRA algorithm and their evolution with the number of agents $n$ in the system. The plot shows (i) the constant theoretical bound on $E_{f}$ (117), in black, (ii) the PEP bounds on the function error of the worst agent $E_{f,\mathrm{worst}}$, scaling sublinearly with $n$, in red and (iii) the PEP bounds for the function error of the average iterate $E_{f}$, in blue. The constant step-size $\alpha^{*}$ has been optimized for this performance criterion and leads to a bound constant with $n$ (plain blue line). Step-sizes can also be optimized with respect to $E_{f,\mathrm{worst}}(15)$, leading to smaller step-sizes $\alpha^{*}_{\mathrm{w}}(n)$ decreasing as $\frac{1}{\sqrt{n}}$. They improve the guarantee for $E_{f,\mathrm{worst}}$ and its scaling with $n$ but they deteriorate the corresponding guarantee for $E_{f}$ (see dashed lines). In this plot, the range of eigenvalues for the averaging matrix is [0.5,0.5][-0.5,0.5] and the local functions are 1-smooth and 0.1-strongly convex.
@@
The maximization of such objectives ensures that agent 1 will be the one with the largest error (on $f$ or $x$). Indeed, if another agent $j$ has a larger error, then the permutation of variables of agents 11 and $j$ in the PEP solution would lead to another feasible PEP solution with a larger objective. Agent 1 will thus receive a specific role in the PEP solution and is not equivalent to the n1n-1 others. We can formulate a compact PEP using two equivalence classes of agents
@@
Firstly, Figure 1 shows that the PEP bound for the average performance $E_{f}$ is well constant over $n$, as predicted by Theorem 7, and outperforms the theoretical bound from [15, Corollary 2], which requires 2750 iterations to guarantee the same error as the PEP bounds after 15 iterations. The theoretical bound is valid for a step-size $\alpha=\frac{1}{4L}$, while the PEP bound is shown for an optimized step-size $\alpha^{*}=\frac{0.78}{L}$, which minimizes the bound. Since the bound is constant over $n$, the value of the optimized step-size can be used for any system size, however, it may depend on other settings, such as the range of eigenvalues for $W$ : $\lambda^{-},\lambda^{+}]$.
There is currently no theoretical guarantee for the worst agent functions error $E_{f,\mathrm{worst}}$ of EXTRA in the literature, but we can analyze it with our compact PEP formulation. Figure 1 shows that the PEP bound for the worst agent $E_{f,\mathrm{worst}}$ (in red) is scaling sublinearly with the number of agents $n$. Using the constant step-size $\alpha^{*}=\frac{0.78}{L}$ (plain red line), we obtain a logarithmic slope of $0.92$. The step-size can also be tuned with respect to $E_{f,\mathrm{worst}}$, and would then have a different value for each value of $n$, denoted $\alpha^{*}_{\mathrm{w}}(n)$. Using $\alpha^{*}_{\mathrm{w}}(n)$ (dashed red line), we obtain a logarithmic slope of $0.67$, which means the scaling with $n$ is less important when using suitable step-sizes. In our case, the step-sizes $\alpha^{*}_{\mathrm{w}}(n)$ are diminishing in $\frac{1}{\sqrt{n}}$. However, these smaller step-sizes deteriorate the average performance $E_{f}$ (dashed blue line).
@@
There is currently no theoretical guarantee for the worst agent iterates error $E_{x,\mathrm{worst}}$ of EXTRA in the literature, but one could derive222using the fact that $\max_{i\in\mathcal{V}}\|x_{i}^{t}-x^{*}\|^{2}\leq\sum_{i=1}^{n}\|x_{i}^{t}-x^{*}\|^{2}$ a conservative bound on $E_{x,\mathrm{worst}}$ from (118), scaling linearly with the number of agents $n$. Using our compact PEP formulation, we can analyze such performance metric accurately and show that the performance of the worst agent $E_{x,\mathrm{worst}}$ for EXTRA actually scales sublinearly with $n$, as shown in Figure 2. Figure 2 shows, in particular, PEP-based bounds for $E_{x}$ and $E_{x,\mathrm{worst}}$. The observations are similar to those of Figure 1. The PEP bound for the average performance $E_{x}$ is constant over $n$, as predicted by Theorem 7 and the corresponding optimized step-size is $\alpha^{*}=\frac{0.78}{L}$. Moreover, the PEP bound for the worst agent $E_{x,\mathrm{worst}}$ is scaling sublinearly with the number of agents $n$. Using the constant step-size $\alpha^{*}=\frac{0.78}{L}$ (plain red line), we obtain a logarithmic slope of $0.82$. As previously, the step-size can also be tuned with respect to $E_{x,\mathrm{worst}}$, and would then have a different value for each value of $n$, denoted $\alpha^{*}_{\mathrm{w}}(n)$. Using the optimal diminishing step-sizes $\alpha^{*}_{\mathrm{w}}(n)$ (dashed red line), we obtain a logarithmic slope of $0.60$, which means the scaling with $n$ is less important when using suitable step-sizes. These smaller step-sizes only slightly deteriorate the average performance $E_{x}$ (dashed blue line).
@@ 6.2 The 80-th percentile of the agent performance
Figure 2: Comparison of the average iterates error $E_{x}$ (119), the worst iterates error $E_{x,\mathrm{worst}}$ (120) and the 80-th percentile iterates error $E_{x,80}$ (121) for $t=15$ iterations of the EXTRA algorithm and their evolution with the number of agents $n$ in the system. The plot shows (i) the PEP bounds on the iterates error of the worst agent $E_{x,\mathrm{worst}}$, scaling sublinearly with $n$, in red, (ii) the PEP bound on the 80-th percentile iterates error $E_{x,80}$, in green, and (iii) the PEP bounds for the average iterates error $E_{x}$, in blue. The constant step-size $\alpha^{*}$ has been hand-tuned for this performance criterion and leads to a bound constant with $n$ (plain blue line). Step-sizes can also be hand-tuned with respect to $E_{f\mathrm{worst}}(15)$, leading to smaller step-sizes $\alpha^{*}_{\mathrm{w}}(n)$. They improve the guarantee for $E_{f,\mathrm{worst}}$ and its scaling with $n$ without deteriorating too much the corresponding guarantee for $E_{f}$ (see dashed lines). In this plot, the range of eigenvalues for the averaging matrix is [0.5,0.5][-0.5,0.5] and the local functions are 1-smooth and 0.1-strongly convex.
In some cases, the performance of the worst agent may not be the most appropriate performance measure. When the network is very large, the performance of the worst agent can be terribly bad, while the average performance is good. Using our PEP approach, we can consider other performance measures that were so far completely out of reach of theoretical analysis. Inspired by statistical approaches, we propose here to analyze the k-th percentile in the distribution of the individual performance of each agent. In this experiment, we have chosen the 80-th percentile, that is, the performance of the worst agent, after excluding the 20% worst agents. In other words, it measures the error at or below which 80% of the agents fall in the worst-case scenario. The performance of the worst agent, analyzed in the previous subsection, can be seen as the 100-th percentile of the agent performance and can be analyzed via PEP using two equivalence classes of agents. When considering smaller percentiles, such as the 80-th, we can formulate a compact PEP using three equivalence classes of agents :
@@
where we assume $n$ can be divided by 5. The class $\mathcal{V}_{1}$ contains the 20% worst agents to exclude, the class $\mathcal{V}_{2}$ contains the agent for which the individual performance will provide the 80-th percentile, and $\mathcal{V}_{3}$ contains all the other agents. The PEP should therefore maximize the performance of agent $i\in\mathcal{V}_{2}$ and impose that the performance of each agent $j\in\mathcal{V}_{j}$ is larger :
$E_{x,80}(t)=\underset{\{f_{A}^{u},G_{A}^{u},G_{B}^{u}G_{E}^{uv}\}_{u,v=1,2,3}}{\mathrm{maximum}}$ $\penalty 10000\ \|x_{i}^{t}-x^{*}\|^{2}\quad\text{ with $i\in\mathcal{V}_{2}$}$ (121)
$\text{ s.t.}\qquad\|x_{j}^{t}-x^{*}\|^{2}$ $\geq\|x_{i}^{t}-x^{*}\|^{2}\qquad\text{for all $j\in\mathcal{V}_{1}$,}$ (122)
All the other $\text{ PEP constraints}.$ (123)
In problem (121), the squared norms $\penalty 10000\ \|x_{i}^{t}-x^{*}\|^{2}$ can be written as $(e_{x^{t}}-e_{x^{*}})^{T}G_{A}^{u}(e_{x^{t}}-e_{x^{*}})$ with $u$ chosen according to the class $\mathcal{V}_{u}$ containing agent $i$. Details about the explicit construction of compact PEP formulations are given in Appendix A.
@@
Figure 3 shows the evolution with k of the k-th percentile of the agent performance for EXTRA, when $n\to\infty$. We see that the limit of the k-th percentile performance is finite for all values of $\mathrm{k}\in(0,100)$ and has a vertical asymptote in $\mathrm{k}=100$. The 100-th percentile computes the performance of the worst agent and grows well to infinity with $n$ (see Figure 2). According to Figure 3, the value of the k-th percentile performance first increases slowly with k, and then it starts to blow up after some threshold on k, depending on $t$, to match the vertical asymptote in $\mathrm{k}=100$. This means that it seems useful to exclude a small part of the worst agents in the performance metric, e.g. 10–20% in the settings we are considering, but not more. Moreover, the k-th percentile of EXTRA for $n\to\infty$ seems to converge geometrically. Indeed, in Figure 3, when considering 5 more iterations in total, the k-th percentile improves by a factor between 2.2 and 3.2, depending on k.
@@
Figures 2 and 3 show the performance of EXTRA for rather well-connected averaging matrices ($\lambda=0.5$). Similar results and observations can be obtained for larger ranges of eigenvalues, i.e. $\lambda\in(0.5,1)$, but this would require more iterations to get small errors.
Figure 3: Evolution with k of the k-th percentile of the agent performance for $t$ iterations of EXTRA, when the number of agents tends to infinity ($n\to\infty$). Different total numbers of iterations $t$ are compared. In this plot, the step-size is $\alpha=0.78$, the range of eigenvalues for the averaging matrix is [0.5,0.5][-0.5,0.5] and the local functions are 1-smooth and 0.1-strongly convex.
@@ 6.3 Performance under local functions heterogeneity
In many applications, the agents hold local functions sharing general properties, e.g. convexity or smoothness, but it is unlikely that all the local functions have exactly the same characterization for these properties. For example, the local functions can be $\mu_{i}$-strongly convex and $L_{i}$-smooth, but with different parameters $\mu_{i}$ and $L_{i}$. In general, the performance is then computed by considering that all the local functions have the worst parameters :
@@
In this experiment, we analyze the gain in the performance guarantee when, instead of considering $f_{i}\in\mathcal{F}_{\mu,L}$ for all $i$, we consider two equivalence classes of agents $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ such that
@@
Figure 4: Evolution with the number of agents $n$ of the worst-case average distance to optimum $E_{x}$ for 15 iterations of EXTRA with two equivalence classes of agents $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ using each a different condition number for their function class. The agents in $\mathcal{V}_{1}$ hold 1-smooth and 0.01-strongly convex local functions ($\kappa_{1}=100$) and the agents in $\mathcal{V}_{2}$ hold 1-smooth and 0.1-strongly convex local functions ($\kappa_{2}=10$). The guarantees are independent of the total number of agents $n$ and only depend on the proportion of agents in each class. In this experiment, the range of eigenvalues for the averaging matrix is [0.5,0.5][-0.5,0.5].
Figure 4 shows the situation where $L=1$, $\mu_{1}=0.01$ and $\mu_{2}=0.1$, for different sizes of classes $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. We observe that the guarantees for EXTRA only depend on the relative composition of each class $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ but not on the total number of agents $n$. This observation has been confirmed by solving the compact PEPs for $n$ tending to $\infty$, as described at the end of Section 5. In Figure 4, when $20\%$ of the agents pass from the class $\mathcal{V}_{1}$ to the class $\mathcal{V}_{2}$, the worst-case guarantee is improved by a constant factor $0.9$. This improving factor can be linked to the worst-case guarantees obtained with a uniform value for $\kappa$ in the local functions ($\kappa=100$ or $\kappa=10$ for all the agents) and motivates the following conjecture :
@@
Let $E_{x}^{\kappa_{1},\kappa_{2}}(t,\theta)$ be the performance guarantee after $t$ iterations of EXTRA with $\theta n$ agents with local functions in $\mathcal{F}_{\mu_{1},L}$ and $(1-\theta)n$ agents with local functions in $\mathcal{F}_{\mu_{2},L}$, for $\theta\in[0,1]$. Let $E_{x}^{\kappa_{1}}(t)$ and $E_{x}^{\kappa_{2}}(t)$ be the worst-case errors of $t$ iterations of the EXTRA algorithm with uniform conditioning for all the local functions, respectively with $\kappa_{1}=\frac{L}{\mu_{1}}$ and $\kappa_{2}=\frac{L}{\mu_{2}}$. By definition, we have
$E_{x}^{\kappa_{1}}(t)=E_{x}^{\kappa_{1},\kappa_{2}}(t,1)\quad\text{ and }E_{x}^{\kappa_{2}}(t)=E_{x}^{\kappa_{1},\kappa_{2}}(t,0)$
@@
$E_{x}^{\kappa_{1},\kappa_{2}}(t,\theta)=\bigl(E_{x}^{\kappa_{1}}(t)\bigr)^{\theta}\bigl(E_{x}^{\kappa_{2}}(t)\bigr)^{1-\theta}\qquad\text{for all $t>0$, $\kappa_{1},\kappa_{2}>0$ and $\theta\in[0,1]$ }.$